OpenStudy (anonymous):

Solve 8(2 1/2x-3/4) - 1/4(1-x)=1/2?? Could somebody please explain these workings 8(5x/2 -3/4) - (1-x)/4=1/2 8((10x-3)/4) -(1-x)/4 = 2/4 (how is it now 2/4???) (what happened to 5x/2) 80x-24-1+x=2 81x=27 x=27/81 x=1/3

OpenStudy (anonymous):

it is hard to read. it the original equation \[8(2\tfrac{1}{2}x-\frac{3}{4})-\frac{1}{4}(1-x)=\frac{1}{2}\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and this is a worked out solution given to you?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok we can follow it as silly as it is

OpenStudy (anonymous):

step one \[2\tfrac{1}{2}=\frac{5}{2}\] right?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

so we get \[8(\frac{5}{2}x-\frac{3}{4})-\frac{1}{4}(1-x)=\frac{1}{2}\]

OpenStudy (anonymous):

right

OpenStudy (anonymous):

step two is to rewrite \[\frac{5}{2}x-\frac{3}{4}=\frac{10x-3}{4}\]apparently. that is just adding up the fractions using a common denominator of 4

OpenStudy (anonymous):

i think maybe that is where you got confused

OpenStudy (anonymous):

where is the 8?

OpenStudy (anonymous):

oh i got it

OpenStudy (anonymous):

ok it is still there, i was just showing one step

OpenStudy (anonymous):

if that is clear, that is where the \[\frac{5}{2}\]went

OpenStudy (anonymous):

now we have \[8(\frac{10x-3}{4})-\frac{1}{4}(1-x)=\frac{1}{2}\]

OpenStudy (anonymous):

as for why is \[\frac{1}{2}=\frac{2}{4}\] it is because they are clearly the same number. multiply top and bottom by 2 and you get it

OpenStudy (anonymous):

and the reason they wrote \[\frac{2}{4}\] instead of \[\frac{1}{2}\] is because now the denominator on the left is the same as the one on the right, so you can just get rid of it.

OpenStudy (anonymous):

that is why the next line should be \[8(10x-3)-(1-x)=2\]

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

don't get attached to this way of solving the problem. you can do it any way that you like. you can pick your own order to distribute, combine like terms etc. this is not the way i would have solved this

OpenStudy (anonymous):

what way would you of solved this???

OpenStudy (anonymous):

after i got rid of the stupid mixed number and saw \[8(\frac{5}{2}x-\frac{3}{4})-\frac{1}{4}(1-x)=\frac{1}{2}\] i probably would have distributed and get \[20x-24-\frac{1}{4}(1-x)=\frac{1}{2}\]

OpenStudy (anonymous):

then * by 4

OpenStudy (anonymous):

then maybe multiply both sides by 4 to clear the fraction. but any way will work

OpenStudy (anonymous):

exactly!

OpenStudy (anonymous):

gotta run. good luck

OpenStudy (anonymous):

thank you